### v1

# See or Say Graphs: Agent-Driven Scalable Graph Understanding with Vision-Language Models

Abstract
[…]
See or Say Graphs: Agent-Driven Scalable Graph Understanding
with Vision-Language Models
Shuo Han1,2*   Yukun Cao1,2*   Zezhong Ding1,2 Zengyi Gao1,2S. Kevin Zhou1,3Xike Xie1,2 1University of Science and Technology of China, China 2Data Darkness Lab, MIRACLE Center, USTC, China 3MIRACLE Center, Suzhou Institute for Advance Research, USTC, China {shuo.han, ykcho, zezhongding, gzy02}@mail.ustc.edu.cn {skevinzhou, xkxie}@ustc.edu.cn
**footnotetext: These authors contributed equally to this work.footnotetext: Corresponding author.
1 Introduction
Recently, vision-language models (VLMs) have shown potential for general-purpose understanding of graph data Tang et al. (2024a); Kong et al. (2025b); Ding et al. (2025b), offering new paradigms for solving real-world problems naturally represented as graphs. A central research focus of this emerging field is to enhance the foundational capability of VLMs to understand graph structures Ren et al. (2024), particularly for large-scale graphs, without relying on external tools such as code execution or software systems Ding et al. (2025a); Feng et al. (2025). To this end, existing studies have explored two primary graph input modalities for VLMs: textual graph descriptions [^1] Many early studies on text-modality-based graph understanding were developed on LLMs. As VLMs extend LLMs with visual capabilities, this paper presents a unified discussion from the perspective of graph understanding with VLMs., which convert structural information into natural language for indirect memory and reasoning over node properties and relations Jiang et al. (2023); Jin et al. (2024); Zhang et al. (2024a); and visual graph representations, which render nodes and edges as structured images to support direct structural perception Li et al. (2024c); Zhu et al. (2025). More recent efforts further combine these modalities to improve structural understanding and complex reasoning on graphs Wei et al. (2024).
Figure 1: Comparison of textual and visual modalities across different types of graph understanding tasks
However, despite these advances, current VLMs still lack a unified framework that systematically addresses the scalability challenges and the difficulties in coordinating textual and visual modalities inherent in graph understanding. Specifically:
Scalability remains a fundamental limitation for both textual and visual graph inputs in VLMs. Given their limited input token capacities, textual graph descriptions are constrained by sequence length, making it difficult to encode the full structure of large graphs Chen et al. (2024a); Yuan et al. (2025a). Visual graph inputs, on the other hand, are restricted by image resolution, resulting in the loss of structural details as graph size increases Zhu et al. (2025).
In addition, VLMs face the lack of an effective mechanism to coordinate different input modalities for graph understanding, which prevents them from fully leveraging the distinct strengths and complementary advantages of textual and visual modalities. As illustrated in Figure 1, under the token length limits of VLMs, the text modality is more effective for simple property tasks such as counting nodes, measuring degrees, and assessing connectivity. In these cases, the information can be directly located and reasoned over with minimal effort. By contrast, the visual modality often performs worse because the limited resolution of global visual graphs blurs structural details, causing VLMs to misidentify relevant elements and leading to performance degradation.
For tasks that require local and structurally complex reasoning, such as shortest-path computation or cycle detection, high-resolution local visual subgraphs offer a clear advantage for VLMs. They provide a “what-you-see-is-what-you-get” view of the topology and naturally support stepwise reasoning. In contrast, textual descriptions are often lengthy and lack structural clarity, leaving VLMs with only indirect language cues and thus little structural grounding for effective complex graph reasoning.
Therefore, a natural conclusion is that realizing the full potential of VLMs in graph understanding requires addressing these two challenges.
This entails, on the one hand, preserving as much task-relevant graph information as possible within limited input tokens, and on the other hand, assigning tasks according to modality strengths: letting the text modality efficiently address simple property tasks, while leveraging the visual modality to directly support local and structurally complex tasks.
To this end, we propose GraphVista, a unified framework for graph understanding with VLMs. For the scalability challenges, GraphVista is inspired by the retrieval-augmented generation (RAG) paradigm Lewis et al. (2020). It organizes graph information hierarchically according to structural importance and stores multi-granularity textual descriptions in a lightweight GraphRAG base. For a given graph understanding task, GraphVista retrieves only the relevant information before input, thereby compressing redundant context and preserving key reasoning elements. When necessary, it can also generate high-resolution visual subgraphs from the retrieved local information to support complex reasoning tasks.
To enable effective coordination across modalities, GraphVista introduces a planning agent that parses graph understanding tasks and routes them to the appropriate modality. The textual modality branch, enhanced with the RAG techniques, efficiently handles simple property tasks by retrieving and reasoning over concise textual descriptions. In contrast, the visual modality branch employs the Visual Graph Thoughts Agent, which constructs a multimodal chain of thought called Visual Graph Thoughts over high-resolution visual subgraphs, enabling VLMs to perform stepwise visual reasoning grounded in explicit structural evidence for structurally complex tasks.
The main contributions of this paper are listed as follows:
  • $\bullet$
    We present the first systematic analysis of the scalability challenges and coordination gaps that hinder VLMs in graph understanding, providing both conceptual insights and empirical evidence.
  • $\bullet$
    We propose GraphVista, the first unified framework for graph understanding with VLMs, designed around two objectives: improving scalability and enabling coordination across modalities. To this end, GraphVista integrates a set of innovative techniques, such as a hierarchical GraphRAG base, a planning agent for task routing, and a Multimodal Graph Thoughts agent for complex reasoning.
  • $\bullet$
    We introduce Grena, a scalable benchmark that supports step-level training of Visual Graph Thoughts and evaluates graph understanding from simple property to complex local tasks.
  • $\bullet$
    Extensive experiments across diverse graph understanding tasks demonstrate the effectiveness of GraphVista, highlighting its ability to scale to large graphs and to exploit the complementary strengths of textual and visual modalities.
2 Related Work
2.1 Text Modality-based Graph Understanding
Most existing studies convert graphs into textual descriptions to enable VLM-based graph understanding, primarily focusing on benchmark construction and capability analysis. Recent benchmarks evaluate VLMs on basic graph element comprehension Wang et al. (2023); Guo et al. (2023); Wu et al. (2024) and graph-theoretic reasoning tasks Chen et al. (2024a); Li et al. (2024b); Tang et al. (2024b); Luo et al. (2024); Yuan et al. (2025a), indicating that the research in this direction remains in an early exploratory phase. Methodologically, Ge et al. (2025); Wang et al. (2025) studied the influence of graph description order, while Cao et al. (2025) identified the “Lost in the Middle” issue in graph sequence. Building upon these findings, Guan et al. (2025) analyzed VLM attention patterns over graph data, and Peng et al. (2025); Zhang et al. (2025) explored task transferability to broader reasoning domains. However, these strategies still face scalability bottlenecks and struggle with complex reasoning tasks.
Other studies address graph problems by leveraging external tools. Li et al. (2024a); Yuan et al. (2025b) employ predefined coding templates for algorithmic reasoning, while Perozzi et al. (2024) trains task-specific GNNs for graph processing. These methods depend on external modules rather than enhancing the intrinsic graph understanding of VLMs, and are thus orthogonal to our work.
2.2 Visual Modality-based Graph Understanding
Visual modality-based graph understanding is still in its infancy. VisionGraph Li et al. (2024c) and VGCure Zhu et al. (2025) benchmark VLMs on graph perception and reasoning, revealing difficulties in capturing structural information. Wei et al. (2024) further introduces a hybrid multimodal representation directly compatible with VLMs. However, these studies remain empirical and fail to address the scalability, structural understanding, and reasoning limitations inherent to vision-only representations.
3 Methodology
Figure 2: Framework of GraphVista
3.1 Overview
In this section, we introduce GraphVista, an agent-driven framework designed to address the challenges of scalability and modality coordination in graph understanding, as shown in Figure 2.
GraphVista operates under the coordination of the planning agent, which orchestrates the entire reasoning workflow. Initially, the planning agent constructs a lightweight Hierarchical GraphRAG Base $\mathcal{K}$ by partitioning the graph based on topological centrality. Upon receiving a question, it parses the task and routes it to the appropriate reasoning modality. For Simple Property Tasks, the text modality-based branch is activated, employing GraphRAG-based retrieval to extract relevant context from $\mathcal{K}$. For Complex Local Tasks that require fine-grained structural reasoning, the task is delegated to the visual modality-based branch. The Visual Graph Thoughts Agent extracts visual subgraphs from $\mathcal{K}$ and performs iterative multimodal reasoning grounded in visual evidence.
The rest of this section details our methodology. We begin by introducing the construction of $\mathcal{K}$ in §3.2. Next, we describe the planning agent-based task routing mechanism in §3.3. Finally, we elaborate on the visual modality-based and text modality-based graph understanding modules in §3.4 and §3.5, respectively.
3.2 Hierarchical GraphRAG Base
A primary challenge for LLMs and VLMs in understanding graph structures is their limited input capacity. This constraint makes it infeasible to encode an entire graph, often causing a loss of critical structural information. To address this limitation, we draw inspiration from RAG and propose building $\mathcal{K}$. Unlike conventional RAG pipelines Guo et al. (2025); Jin et al. (2025) that rely on external knowledge sources, our approach constructs a compact multi-level structural representation directly from the graph and stores it in a lightweight GraphRAG Base, enabling scalable reasoning without external dependencies [^2] For small graphs with $|V|\leq 15$, we skip storage and feed the entire graph directly into the VLMs..
We partition the graph $G=(V,E)$ with $|V|$ nodes into three tiers based on the topological centrality of its nodes. Each tier uses a different storage granularity, significantly reducing storage needs while preserving the structural details essential for downstream tasks.
Tier 1: Core Nodes.
This tier consists of nodes with the highest structural influence, functioning as hubs that govern information flow and shape global graph properties. We employ PageRank, a widely adopted centrality measure that reflects both connectivity and reachability, to identify the top $K_{1}\%$ of nodes as the core nodes. For each core node $v$, its complete 2-hop neighborhood $\mathcal{N}_{2}(v)$ is preserved to capture fine-grained local structural information.
Tier 2: Backbone Nodes.
This tier contains nodes located along intercommunity paths, which are essential for preserving global connectivity. These nodes are selected using betweenness centrality, which quantifies how frequently a node lies on shortest paths between other nodes. The next $K_{2}\%$ of nodes are retained, and for each backbone node $v$, only its 1-hop neighborhood $\mathcal{N}_{1}(v)$ is stored to balance representational fidelity with storage efficiency.
Tier 3: Peripheral Nodes.
The remaining nodes are classified as peripheral, whose structural role is primarily defined through their links to higher-tier nodes. To minimize redundancy, we adopt a conditional storage policy: the 1-hop neighborhood $\mathcal{N}_{1}(v)$ of a peripheral node is stored only if it is not connected to any Tier 1 or Tier 2 node. Otherwise, its structural information is implicitly covered by the neighborhoods of higher-tier nodes, ensuring maximal storage efficiency [^3] In practice, $\mathcal{K}$ can be configured to utilize only a subset of tiers (e.g., only Tier 1) to adapt to storage or computational constraints, providing flexibility across application scenarios..
$\mathcal{K}$ not only addresses scalability but also establishes the information backbone for subsequent multimodal collaborative reasoning. In the text modality-based branch, it enables efficient retrieval of relevant structured context, while in the visual modality-based branch, it guides the precise extraction of local visual subgraphs.
3.3 Planning Agent-based Task Routing
This module employs a planning agent to allocate each task to the optimal reasoning modality, either textual or visual. The routing procedure comprises two stages. In the Task Parsing stage, the agent analyzes the natural language question to extract the task intent and key entities. In the Task Categorization stage, the parsed task is classified and dispatched to the corresponding reasoning module.
Task Parsing
Given a natural language question $Q$ over $G$, the planning agent first performs semantic parsing to identify the task type $T$ and extract the set of key entities $\mathcal{E}$:
$(T,\mathcal{E})=f_{\text{parse}}(Q,G)$ (1)
, where $f_{\text{parse}}$ denotes a semantic parsing function implemented by the planning agent. The agent applies few-shot prompting to infer task semantics, producing a task type $T$ (e.g., neighbor identification) and a set of key entity pairs $\mathcal{E}={(v_{i},v_{j})\mid v_{i},v_{j}\in V}$ referenced in the question. The task type $T$ determines the routing strategy, while the entity set $\mathcal{E}$ provides essential context for downstream modules. To ensure uniform processing, queries involving a single entity (e.g., “What is the degree of node $v_{i}$?”) are reformulated as $(v_{i},\emptyset)$ by introducing a null entity placeholder.
Task Categorization and Execution Planning
Following parsing, the planning agent routes the task to the appropriate module based on $T$ and formulates a high-level execution plan [^4] The planning agent first matches the question against predefined task templates. If no match is found, it falls back to semantic parsing to infer the intent and choose the reasoning modality.. Building on the task definitions established in existing graph understanding benchmarks Chen et al. (2024a); Yuan et al. (2025a); Cao et al. (2025), we propose a task classification framework that leverages the complementary strengths of different graph representation modalities. Specifically, graph tasks are categorized into two classes, each handled by a dedicated module:
Simple Property Task Examples: “What is the total number of nodes in this Graph?”; “Return the degree of node A.” Complex Local Task Examples: “Find the shortest path between nodes A and B.”; “Do nodes A and B share any third-order neighbors?”
Simple Property Tasks.
These tasks involve questioning either macroscopic graph statistics or microscopic, explicitly defined properties. Such information is efficiently and reliably obtained from textual graph descriptions, whereas visual representations often suffer from resolution limitations that obscure fine-grained details. Accordingly, the planning agent routes these tasks to the GraphRAG agent module, providing a high-level retrieval strategy for execution over $\mathcal{K}$.
Complex Local Tasks.
These tasks require reasoning over fine-grained topological structures, such as finding the shortest path between nodes. For such tasks, a visual subgraph offers a concise and intuitive representation that is more effective for complex topological reasoning than verbose textual descriptions. Consequently, the planning agent delegates such tasks to the Visual Graph Thoughts agent, while decomposing the task into algorithm-guided substeps and constructing a high-level reasoning plan.
3.4 Visual Modality-based Graph Understanding
This module leverages visual-modality graph representations to handle complex local problems that require fine-grained topological understanding and multi-step logical inference.
To address the scalability issue of large graphs, we extract a task-relevant $k$-hop subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ with a maximum of $N_{max}$ nodes from $G$ in $\mathcal{K}$, centered on the key entities $\mathcal{E}$ (details are provided in Appendix B). The resulting subgraph is then transformed into a high-resolution visual representation $G_{\text{image}}=f_{\mathrm{viz}}(G^{\prime})$, which preserves the essential topological structure of $G^{\prime}$ and serves as a faithful basis for downstream reasoning.
3.4.1 Visual Graph Thoughts
Existing text-based CoT methods often suffer from pseudo-visual reasoning, in which visual information gradually fades as the reasoning chain lengthens Chen et al. (2024b); Zhang et al. (2024b); Zou et al. (2024); Cheng et al. (2025). To overcome this limitation in multimodal reasoning for complex local tasks, we introduce Visual Graph Thoughts.
After obtaining the visual subgraph $G_{\text{image}}$, the visual reasoning agent $\mathcal{M}_{\text{VRA}}$, instantiated from a VLM, performs iterative multimodal reasoning under the guidance of a high-level plan $\Pi=(\pi_{1},\pi_{2},\dots,\pi_{n})$ generated by the planning agent.
We formalize this process as a state-transition system. At step $t$ ($1\leq t\leq n$), the reasoning state $S_{t}$ is defined by the current visual representation $G_{\text{image}}^{(t-1)}$, the accumulated reasoning history $H_{t-1}$, and the current plan instruction $\pi_{t}$. Conditioned on $S_{t}$, $\mathcal{M}_{\text{VRA}}$ produces an intermediate reasoning output $o_{t}$ and an action $a_{t}$:
$(o_{t},a_{t})=\mathcal{M}_{\text{VRA}}(S_{t})=\mathcal{M}_{\text{VRA}}(G_{\text{image}}^{(t-1)},H_{t-1},{\pi}_{t})$ (2)
If $a_{t}$ invokes $f_{\mathrm{viz}}$ (e.g., node highlighting), the visual representation is updated as $G_{\text{image}}^{(t)}=f_{\mathrm{viz}}(G_{\text{image}}^{(t-1)},a_{t})$; otherwise, $G_{\text{image}}^{(t)}=G_{\text{image}}^{(t-1)}$. This closed-loop process continues until all plan steps are executed.
Finally, $\mathcal{M}_{\text{VRA}}$ aggregates the reasoning trace $H_{n}$ to produce the final answer $A_{\text{final}}=f_{\text{summarize}}(H_{n})$, where $f_{\text{summarize}}$ aggregates intermediate reasoning outputs to generate the final prediction.
3.4.2 Aligning Visual Graph Thoughts with Process-level DPO
To further improve the multimodal reasoning capability of $\mathcal{M}_{\text{VRA}}$ by promoting the generation of reliable visual graph thoughts, we adopt a process-level Direct Preference Optimization (DPO) training strategy Rafailov et al. (2023). Compared to outcome-based methods Kong et al. (2025a); Liu et al. (2025), we model the entire multimodal reasoning trajectory, which provides stable, step-level reinforcement for graph inference tasks.
1. Construction of Process Preference Dataset.
We construct a process preference dataset $\mathcal{D}$, where each sample is a triplet $(\mathbf{x},y_{w},y_{l})$. Here, $y_{w}$ denotes the “Chosen” path representing the preferred reasoning trajectory, while $y_{l}$ denotes the “Rejected” path representing a non-preferred trajectory. The construction details are in Appendix A.
“Chosen” Path ($y_{w}$).
We construct standardized reasoning templates aligned with the logical structure of each graph task. Based on these templates, we generate ground-truth reasoning steps and answers using large-scale synthetic graphs (e.g., ER and BA graphs Li et al. (2005)) and validate their correctness through graph analysis libraries.
“Rejected” Path ($y_{l}$).
To help the VLM distinguish between preferred and non-preferred paths, we construct negative samples $y_{l}$ by modifying $y_{w}$. Specifically, we extend error construction strategies from existing benchmarks Chen et al. (2024a); Yuan et al. (2025a) to multimodal reasoning tasks. The error categories are summarized in Table 5.
2. DPO Training Process.
To align the generation of visual graph thoughts with ground-truth multimodal reasoning trajectories, we fine-tune $\mathcal{M}_{\text{VRA}}$ using DPO on $\mathcal{D}$. For stability, we use a reference policy $\pi_{\text{ref}}$, defined as a copy of $\mathcal{M}_{\text{VRA}}$ that has been supervised to fine-tune in $\mathcal{D}$ and kept fixed during DPO training. The DPO objective minimizes the following loss function:
$\mathcal{L}_{\mathrm{DPO}}=-\mathbb{E}_{(\mathbf{x},y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\!\left(\beta\log\frac{r_{\theta}(y_{w}|\mathbf{x})}{r_{\theta}(y_{l}|\mathbf{x})}\right)\right]$ (3)
, where the policy ratio is defined as $r_{\theta}(y|\mathbf{x})=\pi_{\theta}(y|\mathbf{x})/\pi_{\mathrm{ref}}(y|\mathbf{x})$, and $\beta$ controls the regularization strength. By maximizing the log-probability margin between the preferred ($y_{w}$) and rejected ($y_{l}$) trajectories, the model is guided away from predefined reasoning errors. The resulting policy generates visual graph thoughts that are both logically coherent and tightly grounded in visual graph evidence throughout the inference process.
3.5 Text Modality-based Graph Understanding
For simple property tasks, we adopt GraphRAG-based retrieval for targeted information extraction. Note that existing RAG variants and optimizations Deng et al. (2023); Lee et al. (2025); Gao et al. (2025); Chang et al. (2025) are orthogonal to GraphVista and can be easily integrated to improve retrieval quality. To maintain generality, we employ the standard RAG formulation in this work.
For a given task request $(Q,\mathcal{E})$, we first employs a retrieval function $f_{\text{retrieve}}$ to question $\mathcal{K}$ and extract relevant context, formalized as $C_{\text{retrieved}}=f_{\text{retrieve}}(\mathcal{K},(Q,\mathcal{E}))$. Subsequently, the downstream VLM is prompted with both the original question $Q$ and the retrieved context $C_{\text{retrieved}}$ to generate the final answer.
Table 1: Performance comparison on Grena and GraphSQA Benchmark. Red and Blue highlight the best and second-best results respectively.
Method Model Grena GraphSQA
Simple Complex Overall Simple Complex Overall
Text-only InternVL3-9B 0.1211 0.0287 0.0829 0.4788 0.2498 0.3169
GLM-4.1V-9B 0.0157 0.0071 0.0121 0.2532 0.1670 0.1929
Qwen2.5-VL-7B 0.0096 0.0095 0.0096 0.2340 0.1122 0.1488
Qwen3-8B 0.0065 0.0056 0.0061 0.1351 0.1352 0.1352
Gemma-3-12B 0.0369 0.0358 0.0364 0.5917 0.3245 0.4033
GraphPRM (DPO) Qwen2.5-VL-7B (DPO) 0.0332 0.0087 0.0231 0.7198 0.3455 0.4550
Qwen3-8B (DPO) 0.0060 0.0049 0.0055 0.3629 0.2357 0.2728
Gemma-3-12B (DPO) 0.0417 0.0335 0.0383 0.6235 0.3455 0.4195
GraphInsight InternVL3-9B 0.1788 0.0056 0.1073 0.7564 0.2315 0.3860
GLM-4.1V-9B 0.1849 0.0016 0.1092 0.4313 0.3273 0.3583
Qwen2.5-VL-7B 0.1925 0.0085 0.1165 0.5769 0.2293 0.3348
Gemma-3-12B 0.2919 0.0236 0.1811 0.7372 0.2478 0.3779
GraphToken Qwen2.5-VL-7B 0.0981 0.0782 0.0899 0.0388 0.1770 0.1376
Qwen3-8B 0.1589 0.0757 0.1245 0.1742 0.1508 0.1579
Gemma-3-12B 0.0158 0.0576 0.0331 0.0566 0.1091 0.0927
GITA InternVL3-9B 0.1751 0.1098 0.1481 0.3910 0.1755 0.2402
GLM-4.1V-9B 0.1651 0.1175 0.1454 0.3397 0.1562 0.2112
Qwen2.5-VL-7B 0.1351 0.0632 0.1054 0.3205 0.1596 0.2079
Gemma-3-12B 0.1732 0.1435 0.1609 0.2692 0.1739 0.2026
GraphVista (Ours) InternVL3-9B 0.9357 0.3578 0.6970 0.7970 0.6886 0.7211
GLM-4.1V-9B 0.9387 0.2727 0.6637 0.7954 0.6657 0.7046
Gemma-3-12B 0.9362 0.3477 0.6932 0.7991 0.6815 0.7168
Gemma-3-4B (DPO) 0.9362 0.3397 0.6899 0.7943 0.6780 0.7129
Qwen2.5-VL-7B (DPO) 0.9367 0.3912 0.7114 0.7973 0.6806 0.7156
4 Evaluation
4.1 Experimental Setup
4.1.1 Baselines
We compare GraphVista with representative state-of-the-art baselines across three categories: (a) Text-based methods, which reason solely over textual graph descriptions, including the standard Text-only approach, GraphPRM Peng et al. (2025) trained via DPO, and GraphInsight Cao et al. (2025). (b) GNN-based methods, which encode structural information into latent representations using graph neural networks. A representative example is GraphToken [^5] These methods are orthogonal to GraphVista, focusing on graph encoding rather than improving the intrinsic graph understanding of VLMs. We include them for completeness. Perozzi et al. (2024). (c) Hybrid methods, exemplified by GITA Wei et al. (2024), which integrate both visual and textual graph modalities into VLMs.
4.1.2 Benchmarks and Evaluation Tasks
Existing graph understanding benchmarks suffer from limited coverage of graph scales and topologies, a lack of task categorization grounded in fundamental graph properties, and insufficient support for evaluating multi-step visual reasoning. To address these gaps, we introduce Grena, a large-scale, multi-task benchmark we designed to evaluate graph understanding of LLMs and VLMs, which encompasses a wide range of task types found in existing benchmarks Wang et al. (2023); Yuan et al. (2025a); Zhu et al. (2025).
Grena spans a wide diversity of graph scales, with node counts ranging up to 2050, and encompasses 18 distinct task types, which are organized into two categories: (1) 7 Simple Property Tasks (Simple) focus on querying basic properties and local structures, and (2) 11 Complex Local Tasks (Complex) require fine-grained topological information and performing fine-grained inference. Further details are provided in Appendix A.
In our experiments, we use Grena to evaluate graph understanding performance on large-scale graphs ($|V|\in[50,2050]$), covering a total of 13,860 tasks, which are split into 8,136 simple and 5,724 complex instances. Complementarily, we use the GraphSQA benchmark Cao et al. (2025) for small-graph understanding, with $|V|$ in the range of $[15,50]$. Across all tasks and benchmarks, we use Accuracy (ACC) as the evaluation metric. Methods that require DPO are trained on $\mathcal{D}$.
(a) Complex (Gemma-3)
(b) Simple (Gemma-3)
(c) Complex (Qwen2.5-VL)
(d) Simple (Qwen2.5-VL)
Figure 3: Analysis on Graphs with Different $|V|$
Table 2: Ablation study on reasoning strategies of GraphVista. The best results within each VLM group are bolded, and the second-best results are underlined.
Strategy Simple Complex Overall
InternVL3-9B
CoT (Baseline) 0.9368 0.1422 0.6086
Visual Graph Thoughts (Text) 0.2220 0.6416
Visual Graph Thoughts (Visual) 0.3578 0.6977
Gemma-3-12B
CoT (Baseline) 0.9362 0.1406 0.6076
Visual Graph Thoughts (Text) 0.2447 0.6506
Visual Graph Thoughts (Visual) 0.3477 0.6932
GLM-4.1V-9B
CoT (Baseline) 0.9387 0.1406 0.6091
Visual Graph Thoughts (Text) 0.2300 0.6460
Visual Graph Thoughts (Visual) 0.2727 0.6637
Qwen2.5-VL-7B (DPO)
CoT (Baseline) 0.9367 0.1414 0.6083
Visual Graph Thoughts (Text) 0.3429 0.6915
Visual Graph Thoughts (Visual) 0.3912 0.7114
4.1.3 Models
We instantiate GraphVista and all baselines using recent open-source, competitive VLMs, including Gemma-3-12B, Gemma-3-4B Kamath et al. (2025), Qwen2.5-VL-7B Bai et al. (2025), InternVL3-9B Zhu and et al (2025), and GLM-4.1V-9B-Thinking Hong et al. (2025), a reasoning-oriented VLM. For text-based methods, we include more capable LLMs such as Qwen3-8BYang et al. (2025) to ensure a fairer comparison.
4.2 Performance
As shown in Table 1, GraphVista significantly outperforms all baseline methods on both Grena and GraphSQA benchmarks.
4.2.1 Simple Property Tasks
[…]
4.2.2 Complex Local Tasks
[…]
Table 3: Ablation results comparing frozen base models (“Frozen”) with models further optimized via DPO (“DPO”).Red and Blue highlight the best and second-best results respectively.
Model Strategy Simple Complex Overall
Gemma-3-4B Frozen 0.9362 0.2933 0.6707
DPO 0.3397 0.6899
Qwen2.5-VL-7B Frozen 0.9367 0.3240 0.6837
DPO 0.3912 0.7114
4.3 Ablation Study
4.3.1 Effectiveness of Reasoning Strategies
We evaluate the effect of visual grounding by comparing Visual Graph Thoughts (Visual) with its text-only variant and the standard textual CoT, where visual subgraphs are replaced by textual descriptions. In Table 2, the visual variant consistently outperforms the baselines, especially on complex tasks, achieving up to $1.1\times$ and $2.8\times$ higher accuracy than the text-only and CoT settings, respectively. These highlight the importance of explicit structural grounding, as purely textual reasoning often yields ambiguous interpretations.
4.3.2 Impact of DPO
We evaluate the impact of process-level DPO by comparing fine-tuned models with their frozen-base counterparts. As reported in Table 3, DPO enhances performance across all settings, with the most significant gains on complex tasks. For instance, DPO boosts the accuracy of Qwen2.5-VL-7B on complex tasks from 0.3240 to 0.3912. These results validate that aligning model generation with expert visual reasoning trajectories is highly effective for complex local tasks.
(a) Gemma-3 (b) Qwen2.5-VL
Figure 4: Hyperparameter analysis on $k$ and $N_{max}$
Figure 5: Hyperparameter analysis on storage strategy of $\mathcal{K}$
4.4 Hyperparameter Analysis
4.4.1 Subgraph Extraction Parameters $k$ and $N_{max}$
We analyze the impact of $k$ and $N_{max}$ on performance for complex local tasks. As shown in Figure 4, Gemma-3 reaches its highest accuracy with a 2-hop neighborhood, indicating that it can leverage a broader structural context. In contrast, Qwen2.5-VL performs best with a 1-hop neighborhood, suggesting that larger subgraphs may introduce extraneous information. These observations point to the need for adapting subgraph extraction strategies to the characteristics of different VLM architectures.
4.4.2 Storage Strategy of $\mathcal{K}$
We analyze the effect of pruning node tiers in $\mathcal{K}$, which are defined by the parameters $K_{1}$% and $K_{2}$%. As demonstrated in Figure 5, performance on simple property tasks improves as more tiers of the graph are stored. Retaining just the top $K_{1}$% of nodes (Tier 1) along with their 2-hop neighborhoods already provides a strong baseline for localized retrieval. Adding the next $K_{2}$% of nodes (Tier 2) enhances performance further, although the gains diminish when including peripheral nodes (Tier 3). While full graph storage is essential for precise macroscopic tasks where any pruning leads to information loss, these findings reveal a clear trade-off between performance and storage efficiency for general reasoning tasks.
5 Conclusion
In this paper, we propose GraphVista, a unified framework addressing scalability and modality coordination in graph understanding. To improve scalability, GraphVista uses a lightweight hierarchical GraphRAG base to compress and store task-relevant information. For modality coordination, the planning agent routes tasks to the most suitable modality, text modality for simple property tasks or visual modality for complex local tasks. Future work will extend GraphVista to more complex graph types with labels and semantics.
Limitations
The proposed GraphVista framework offers a promising approach for scalable, modality-coordinated graph understanding, while also presenting several potential extensions that could inspire future research directions.
First, while the framework currently focuses on graph topology, its modular design is intended to facilitate extensions for handling rich semantic information. For instance, a semantic extraction module could be easily incorporated to process node attributes and edge labels. This would broaden the framework’s applicability to domains such as knowledge graph reasoning and question answering over heterogeneous networks. Additionally, for the lightweight GraphRAG base, we adopted a general and efficient implementation to ensure the framework’s versatility. We note that many advanced optimization strategies, such as sophisticated indexing and retrieval techniques, are orthogonal to our core contributions. A key advantage of GraphVista’s modular design is that these specialized modules can be readily integrated into our framework. Therefore, future work could easily incorporate state-of-the-art components to further enhance the performance of our framework, demonstrating its flexibility.
Appendix A Grena Benchmark
Existing graph understanding benchmarks suffer from three major limitations: (i) limited coverage of graph scales and topological structures; (ii) lack of task categorization grounded in fundamental graph properties; and (iii) insufficient support for modeling and evaluating multi-step visual reasoning in graph inference.
To address these gaps, we propose the Grena Benchmark, a large-scale, multi-task benchmark with step-level visual reasoning capability. It is designed to rigorously evaluate and advance the cognitive and reasoning abilities of Vision-language models (VLMs) in graph domains.
The core design of Grena Benchmark is reflected in the following three aspects:
Table 5: Categories of errors used to construct $y_{l}$.
Textual and Logical Errors
Factual Errors Incorrect identification of a neighbor set $\mathcal{N}(v)$ or an edge weight $w(e)$.
Logical Errors Faults in algorithmic reasoning (e.g., using edge count instead of weight in a shortest-path task).
Computation Errors Mistakes in numerical operations such as counting.
Omitted Steps Skipping essential reasoning steps.
Multimodal Errors
Element Misrecognition Misidentifying nodes or edges in the visual representation.
Visual Neglect Failure to incorporate visual information, with reasoning based only on text.
Text–Visual Inconsistency Mismatch between textual reasoning and visual grounding.
Visualization Misuse Incorrect or missing calls to visualization functions.
A.1 Scalability and Structural Diversity
A primary objective of Grena is to systematically evaluate the scalability graph understanding capabilities of VLMs. To this end, the benchmark includes graphs that span a wide range of scales, with node counts from 50 to 2,050. This diversity in size is essential for assessing how model performance varies with increasing graph size and structural complexity.
In addition to scale, Grena incorporates diverse topological structures by including both Erdős–Rényi (ER) random graphs and Barabási–Albert (BA) scale-free networks. While ER graphs provide a baseline for performance on uniform structures, BA graphs feature power-law degree distributions and hubs, closely resembling the topology of many real-world systems. This combination of scales and topologies creates a robust testbed for evaluating model generalization across heterogeneous graph structures.
A.2 Categorization of Graph Understanding Tasks
To provide a structured assessment of reasoning abilities, Grena Benchmark organizes tasks into two categories:
  • Simple QA Tasks. These tasks focus on querying basic properties and local structures of graphs, such as “What is the total number of nodes?” or “What is the degree of node $v_{A}$?” They primarily evaluate a model’s ability to store and retrieve graph information. With a highly regularized form, these tasks emphasize breadth of coverage and scalability.
  • Complex Reasoning Tasks. These tasks require models to integrate fine-grained topological structures and perform logical reasoning. For example, shortest-path problems demand identifying the optimal path from node $v_{A}$ to node $v_{B}$. Compared with retrieval tasks, these tasks probe deeper reasoning capabilities by requiring structural integration and inference.
A.3 Process-Level Supervision for Visual Reasoning
A key contribution of the Grena Benchmark is its native support for training models on step-level visual reasoning. For tasks that require a chain-of-thought process (e.g., pathfinding), the benchmark provides not only the final answers but also intermediate visual exemplars. These include the original graph and a series of "state graphs" that progressively highlight key nodes, edges, or subgraphs. This design provides dense supervisory signals, making it suitable for both Supervised Fine-Tuning (SFT) and alignment techniques like Direct Preference Optimization (DPO), thereby enabling VLMs to learn explicit reasoning processes.
To facilitate these training paradigms, we construct a high-quality, process-centric dataset where each sample is a triplet $(\mathbf{x},y_{w},y_{l})$, representing the input, a preferred reasoning path (Chosen), and a non-preferred one (Rejected):
  • Supervised Fine-Tuning Data (The "Chosen" Path, $y_{w}$). The "Chosen" paths serve as gold-standard examples for SFT. They are generated based on human-annotated, standardized reasoning templates that align with the logical flow of each task. We use large-scale synthetic graphs (e.g., Erdős–Rényi and Barabási–Albert models), and programmatically validate the correctness and completeness of all reasoning steps using libraries like NetworkX.
  • Direct Preference Optimization Data (The "Rejected" Path, $y_{l}$). To create preference pairs $(y_{w},y_{l})$ for DPO, the "Rejected" paths are systematically generated by introducing errors into the "Chosen" paths. These error-construction strategies are adapted from established textual reasoning benchmarks and extended to multimodal contexts. The resulting paths feature common reasoning pitfalls (as detailed in Table 5), teaching the model to distinguish between correct and flawed reasoning.
Table 6: Task types in the Grena Benchmark: definition, expected output and category.
Task Name Definition and Expected Output
Bipartite Detection Determine whether $G=(V,E)$ is bipartite (2-colorable). Output: boolean (is bipartite?), optionally a bipartition $(S,T)$.
Clique Detection Detect the existence of (or find) a clique of specified size $k$, or return maximal cliques. Output: boolean (exists $k$-clique?) or node set(s) forming a clique.
Common Third-Order Neighbors Given nodes $u,v$, compute the intersection $N^{(3)}(u)\cap N^{(3)}(v)$ where $N^{(3)}(\cdot)$ denotes nodes at distance exactly 3. Output: set of nodes.
Connectivity Check Decide whether $G$ is connected (single connected component). Output: boolean.
Connectivity Analysis Full analysis of connectivity: number of connected components, sizes of components, and component ID/size for a query node. Output: integer(s) and list(s) (component counts and sizes).
Critical Node Detection Identify articulation points (nodes whose removal increases the number of connected components) or nodes critical to connectivity. Output: set of node ids.
Cycle Detection Determine whether any cycle exists; optionally return an example cycle (sequence of nodes). Output: boolean and, if requested, a cycle path.
Edge Count Return the total number of edges $|E|$. Output: integer.
Edge Existence Given a pair $(u,v)$, decide whether edge $(u,v)\in E$. Output: boolean.
Find Connected Edges Given a node $v$ (or a small subgraph), return the set of incident edges or the edges that connect the queried substructure. Output: set of edges (pairs).
Highest-Degree Neighbor For a query node $v$, return the neighbor $w\in N(v)$ with the largest degree $\deg(w)$ (tie-breaking rule: smallest id or deterministic rule). Output: node id.
Neighbor Connections Measure connections among the neighbors of $v$: e.g., number of edges in the induced subgraph $G[N(v)]$. Output: integer (edge count) or list of edges.
Node Count Return the total number of nodes $|V|$. Output: integer.
Node Degree Given node $v$, return its degree $\deg(v)$ (or weighted degree if weights provided). Output: integer (or numeric).
Planarity Testing Determine whether $G$ is planar; optionally, produce a planar embedding if planar. Output: boolean and (optional) embedding.
Shortest Path Given nodes $u,v$, compute a shortest path (and its length). Supports unweighted or weighted variants. Output: sequence of nodes (path) and length.
Third-Order Neighbors For node $v$, return $N^{(3)}(v)$, the set of nodes at graph distance exactly 3 from $v$. Output: set of nodes.
Triangle Counting Count all triangles (3-cycles) in $G$, and/or enumerate them. Output: integer (triangle count) or list of 3-tuples (node triples).
Appendix B Subgraph Extraction
To mitigate the scalability challenges of large graphs, we first extract a task-relevant local subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ from $G$ stored in $\mathcal{K}$, based on the key entities $\mathcal{E}$, formalized as:
$G^{\prime}=f_{\text{extract}}(G,\mathcal{E})$ (4)
, where the extraction strategy $f_{\text{extract}}$ is determined by the structural characteristics of $\mathcal{E}$, and can be categorized as follows:
Ego-centric
For problems involving a single entity, where $\mathcal{E}=(v_{i},\emptyset)$, we extract its $k$-hop neighborhood subgraph $G^{\prime}_{v_{i},k}=(V^{\prime},E^{\prime})$. The node set $V^{\prime}$ is defined as:
$V^{\prime}=\{v\in V\mid\text{dist}_{G}(v,v_{i})\leq k\}$ (5)
, where $\text{dist}_{G}(u,v)$ denotes the shortest path distance between nodes $u$ and $v$ in $G$. The edge set $E^{\prime}$ is induced from $G$ by $V^{\prime}$.
Multi-centric
For problems concerning relationships between multiple entities, where $\mathcal{E}=(v_{s},v_{t})$, we extract paths connecting the source $v_{s}$ and target $v_{t}$. We employ Yen’s K-shortest paths to obtain $K$ candidate paths $\mathcal{P}=\{P_{1},\dots,P_{K}\}$. We then construct a set $V_{\mathcal{P}}=\bigcup_{j=1}^{K}V(P_{j})$ of all nodes on these paths. Finally, a 1-hop neighborhood expansion is performed on all nodes in $V_{\mathcal{P}}$ to form the final subgraph node set $V^{\prime}=\bigcup_{v\in V_{\mathcal{P}}}\{u\in V\mid\text{dist}_{G}(u,v)\leq 1\}$. The edge set $E^{\prime}$ is induced by $V^{\prime}$ from $G$.
Appendix C Subgraph Pruning Algorithms
To maintain visual clarity and adhere to computational constraints, particularly for large graphs, the size of an extracted subgraph $G^{\prime}$ is bounded by a maximum node count, $N_{max}$. In instances where the initial extracted subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ exceeds this threshold, a principled pruning procedure is invoked to reduce its size while preserving the most topologically significant nodes. The specific pruning strategy is contingent upon the extraction method used, namely ego-centric or multi-centric, as detailed in Algorithm 1 and Algorithm 2, respectively.
C.1 Ego-centric Subgraph Pruning
This approach, detailed in Algorithm 1, addresses subgraphs generated around a single central node, $v_{i}$. The core principle is to retain nodes most relevant to this central entity. All nodes except for $v_{i}$ are considered candidates for removal and are sorted into a priority list for pruning based on the hierarchical key $(-d(v,v_{i}),t(v),c(v))$. This multi-level sorting criterion prioritizes the removal of nodes based on the following order:
  1. 1.
    Distance: Nodes farthest from the central node $v_{i}$ (descending distance $d$) are removed first.
  2. 2.
    Structural Tier: Among nodes at the same distance, those in lower-importance tiers within the Hierarchical GraphRAG Base $\mathcal{K}$ (ascending tier $t(v)$) are prioritized for removal.
  3. 3.
    Centrality: For nodes with the same distance and tier, those with a lower centrality score (ascending centrality $c(v)$) are removed.
By systematically applying these criteria, the algorithm ensures that the pruned subgraph retains the most structurally-important neighborhood around the central node.
C.2 Multi-centric Subgraph Pruning
This strategy, presented in Algorithm 2, is designed for subgraphs constructed around a set of K-shortest paths, $V_{\mathcal{P}}$, connecting multiple entities. The primary objective is to preserve the integrity of these critical paths. Therefore, all nodes lying on these paths ($V_{\mathcal{P}}$) are explicitly protected from pruning. Candidate nodes for removal are those not on the primary paths, and their sorting priority is determined by the key $(t(v),c(v))$, which prioritizes removal based on:
  1. 1.
    Structural Tier: Nodes in lower-importance tiers (ascending tier $t(v)$).
  2. 2.
    Centrality: Among nodes in the same tier, those with a lower centrality score (ascending centrality $c(v)$).
This approach effectively reduces the subgraph’s complexity while guaranteeing that the core relational structure between the key entities remains intact.
Input : $G^{\prime}=(V^{\prime},E^{\prime})$: Initial $k$-hop neighborhood subgraph.
$v_{i}$: Central node for extraction.
$N_{\max}$: Maximum allowed number of nodes.
$\mathcal{K}$: Hierarchical GraphRAG Base.
Output : $G^{\prime}_{\text{pruned}}$: Pruned subgraph.
1ex
if $|V^{\prime}|\leq N_{\max}$ then return $G^{\prime}$
$V_{\text{cand}}\leftarrow V^{\prime}\setminus\{v_{i}\}$
// Sort by priority: distance $\downarrow$, tier $\uparrow$, centrality $\uparrow$
Sort $V_{\text{cand}}$ into list $L$ by key $(-d(v,v_{i}),t(v),c(v))$
$n_{\text{prune}}\leftarrow|V^{\prime}|-N_{\max}$
$V_{\text{prune}}\leftarrow L[1:n_{\text{prune}}]$
$V_{\text{kept}}\leftarrow V^{\prime}\setminus V_{\text{prune}}$
$G^{\prime}_{\text{pruned}}\leftarrow$ subgraph of $G^{\prime}$ induced by $V_{\text{kept}}$
return $G^{\prime}_{\text{pruned}}$
Algorithm 1 Ego-centric Subgraph Pruning
Input : $G^{\prime}=(V^{\prime},E^{\prime})$: Subgraph expanded from $K$-shortest paths.
$V_{\mathcal{P}}$: Nodes on the $K$-shortest paths.
$N_{\max}$: Maximum allowed number of nodes.
$\mathcal{K}$: Hierarchical GraphRAG Base.
Output : $G^{\prime}_{\text{pruned}}$: Pruned subgraph.
1ex
if $|V^{\prime}|\leq N_{\max}$ then return $G^{\prime}$
$V_{\text{cand}}\leftarrow V^{\prime}\setminus V_{\mathcal{P}}$
// Sort by priority: tier $\uparrow$, centrality $\uparrow$
Sort $V_{\text{cand}}$ into list $L$ by key $(t(v),c(v))$
$n_{\text{prune}}\leftarrow|V^{\prime}|-N_{\max}$
$V_{\text{prune}}\leftarrow L[1:n_{\text{prune}}]$
$V_{\text{kept}}\leftarrow V^{\prime}\setminus V_{\text{prune}}$
$G^{\prime}_{\text{pruned}}\leftarrow$ subgraph of $G^{\prime}$ induced by $V_{\text{kept}}$
return $G^{\prime}_{\text{pruned}}$
Algorithm 2 Multi-centric Subgraph Pruning
Table 7: Key Hyperparameters and Framework Configurations.
Category Parameter Value
Model & Inference Configuration
Model Path MODEL_PATH (Path to VLM)
Tensor Parallel TENSOR_PARALLEL_SIZE 1
GPU Memory Util. GPU_MEMORY_UTILIZATION 0.4
Max Model Length MAX_MODEL_LEN 4096
Generation Sampling Parameters
Temperature TEMPERATURE 0.01
Max Tokens MAX_TOKENS 2048
Top P TOP_P 0.9
Subgraph Extraction
Max Subgraph Nodes max_nodes 25
Max Hops max_hops 2
Appendix D Notations
This section summarizes all notations used throughout this paper, as in Table 8.
Table 8: Summary of notations used in this paper.
Notation Description
$G=(V,E)$ Graph with node set $V$ and edge set $E$.
$\mathcal{K}$ The Hierarchical GraphRAG Base.
$Q,T,\mathcal{E}$ NL question, its task type, and key entities.
$f_{\text{parse}}$ Semantic parsing function for questions.
$\mathcal{N}_{k}(v)$ $k$-hop neighborhood of node $v$.
$|V|,N_{\text{max}}$ Total nodes in graph; max nodes in a subgraph.
$K_{1}\%,K_{2}\%$ Node percentage in Tier 1 (Core) and Tier 2 (Backbone).
$G^{\prime}=(V^{\prime},E^{\prime})$ A subgraph extracted from the original graph $G$.
$f_{\text{viz}},G_{\text{image}}$ Subgraph visualization function and its visual output.
$\mathcal{M}_{\text{VRA}}$ The Visual Reasoning Agent.
$\Pi,S_{t},H_{t}$ Reasoning plan, state at step $t$, and history up to $t$.
$(o_{t},a_{t})$ Intermediate output and action at step $t$.
$C_{\text{retrieved}}$ Retrieved context from the GraphRAG Base $\mathcal{K}$.
$f_{\text{retrieve}}$ Retrieval function for querying $\mathcal{K}$.
$\mathcal{D},(x,y_{w},y_{l})$ Preference dataset and sample (input, chosen, rejected).
$\pi_{\theta},\pi_{\text{ref}}$ DPO policy and fixed reference models.
$r_{\theta}(y|x)$ Ratio of policy probabilities $\pi_{\theta}(y|x)/\pi_{\text{ref}}(y|x)$.
$\beta,\mathcal{L}_{\text{DPO}}$ DPO regularization strength and its loss function.
$f_{\text{extract}}$ The subgraph extraction function.
$dist_{G}(u,v)$ Shortest path distance between nodes $u$ and $v$.
$V_{\mathcal{P}}$ Set of nodes on the K-shortest paths.
$d(v,v_{i}),t(v),c(v)$ Node’s distance to center $v_{i}$, structural tier, and centrality.
Appendix E Experimental Settings
Our experiments are conducted using 8 NVIDIA H20 GPUs. The framework is implemented in Python 3.11 with CUDA Version 12.6, leveraging the vLLM library for efficient inference of the VLMs.
Appendix F Implementation Details
Key hyperparameters and framework configurations are detailed in Table 7.
(a) Original graph
(b) Identify neighbors
(c) Check edges
(d) Triangle confirmed
Figure 6: Visualizing the process of Triangle Detection for node A. The algorithm identifies node A’s neighbors and then checks for edges between them to confirm a 3-cycle.
(a) Identify start (B) and end (F)
(b) Explore from node B
(c) Path extended to node D
(d) Path extended to node E
(e) Shortest path found
Figure 7: Visualizing a breadth-first search for the Shortest Path from node B to F. The algorithm explores neighbors layer-by-layer until reaching the destination.
(a) Step 1: Identify target node A
(b) Step 2: Highlight direct neighbors
Figure 8: Visualizing Neighbor Retrieval for a target node A. The process identifies all nodes that are directly connected to node A.
Appendix G Example of Visual Graph Thought Chains
This section provides concrete examples of the step-by-step "Visual Graph Thoughts" process, a core component of our framework for solving complex local reasoning tasks. Each figure-set deconstructs a graph-based problem into a sequence of visual states. This methodology allows the VLM to perform grounded reasoning by identifying and highlighting relevant nodes and edges at each step, culminating in a final solution.
Figure 6 illustrates the process for Triangle Detection. The chain begins by identifying the neighbors of a central node, then visually inspects for edges between those neighbors to confirm a 3-cycle. For Shortest Path finding, shown in Figure 7, the model emulates a breadth-first search, extending the path layer-by-layer from the source until the destination is reached. Finally, Figure 8 depicts the fundamental operation of Neighbor Retrieval, where the model identifies and highlights all nodes directly adjacent to a given target. These examples demonstrate how Visual Graph Thoughts provide explicit, verifiable reasoning grounded in visual evidence.
Appendix H Prompt and QA Templates
This section provides prompt templates for evaluating the graph understanding capabilities of Vision-Language Models (VLMs) within the GraphVista framework. The templates are categorized into Simple Property Tasks and Complex Local Tasks, reflecting the task routing mechanism of the planning agent as described in Section 3.3 of the main paper. Each template includes an example prompt, the expected model behavior based on the GraphVista architecture, and the rationale behind the chosen modality.
H.1 Simple Property Tasks: Text Modality with GraphRAG
For simple property tasks, the planning agent routes the query to the text-modality branch. This branch utilizes GraphRAG-based retrieval on the Hierarchical GraphRAG Base ($\mathcal{K}$) to efficiently extract factual information without needing visual inference.
H.2 Complex Local Tasks: Visual Modality with Visual Graph Thoughts
For tasks requiring nuanced topological understanding or multi-step reasoning, the planning agent delegates the task to the visual modality branch. This module extracts a relevant subgraph, visualizes it, and employs the Visual Graph Thoughts agent to perform grounded, step-by-step multimodal reasoning.
Task 1: Node Existence and Properties Prompt Template: "Determine if node <node_id> exists in the graph. If so, what is its degree?" GraphVista Behavior: The planning agent categorizes this as a simple property task. The model performs a targeted retrieval query against the GraphRAG Base $\mathcal{K}$ to find the entry for <node_id>. The retrieved context, containing the node’s 1-hop neighborhood information, is then used to generate a direct answer. Rationale: This task involves retrieving explicitly stored attributes. The textual representation in the GraphRAG base is optimized for such direct lookups, making it efficient and reliable.
Task 2: Edge Existence and Weight Prompt Template: "Is there a direct edge between node <node_id_1> and node <node_id_2>? If yes, provide its weight." GraphVista Behavior: The query is identified as a simple property task. The system retrieves the adjacency information for <node_id_1> from $\mathcal{K}$. The VLM then processes this structured text to check for a connection to <node_id_2> and extracts the corresponding weight attribute. Rationale: Like node property lookups, edge verification relies on retrieving specific, pre-processed information. The textual modality avoids the potential for visual ambiguity (e.g., overlapping edges or nodes) that could occur in a complex graph visualization.
Task 3: Shortest Path Identification Prompt Template: "Find and list the sequence of nodes that form the shortest path between node <node_id_1> and node <node_id_2>." GraphVista Behavior: The planning agent categorizes this as a complex local task. The Visual Graph Thoughts Agent extracts a multi-centric subgraph centered around the shortest path candidates between <node_id_1> and <node_id_2>. This high-resolution subgraph is visualized. The VLM then initiates a Visual Graph Thought process, highlighting nodes and edges step-by-step in the image to trace the optimal path, verbalizing its reasoning at each step before presenting the final sequence. Rationale: Textual descriptions of paths can be convoluted and difficult for VLMs to reason over effectively. A visual representation provides a “what-you-see-is-what-you-get” view of the topology, which is more intuitive for complex structural reasoning and naturally supports stepwise inference grounded in visual evidence.
Task 4: Triangle Detection Prompt Template: "Does a 3-cycle (triangle) involving node <node_id> exist? If so, list the nodes of one such triangle." GraphVista Behavior: This is routed to the visual modality. An ego-centric subgraph around <node_id> is extracted from $\mathcal{K}$ and visualized. The Visual Graph Thoughts agent first identifies the neighbors of <node_id> in the image. It then visually inspects for edges connecting any two of these neighbors. If such an edge is found, the VLM confirms the existence of a triangle and outputs the three nodes forming the cycle. Rationale: Detecting local structural patterns like cycles is inherently a topological problem. Visual inspection allows the VLM to directly perceive the connectivity pattern, whereas a text-based approach would require inefficient and error-prone traversal of adjacency lists to check for path closures.